# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v3

@@ 7 Conclusion
In this work, we introduced PnP-CoSMo, a modular approach to multi-contrast reconstruction which combines content/style modeling with iterative reconstruction, offering the reconstruction quality of state-of-the-art end-to-end methods with improved generalizability despite lenient training data requirements. At its core is the content consistency operation, which provides regularization at the semantic level of the image content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-CoSMo provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-CoSMo produced diagnostic-quality images at =3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving and extending the content/style model, reducing reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.
@@
Chinmay Rao: Conceptualization, Methodology, Data curation, Software, Investigation, Validation, Formal analysis, Visualization, Writing – original draft. Matthias van Osch: Funding acquisition, Supervision, Conceptualization, Validation, Writing – review & editing. Nicola Pezzotti: Funding acquisition, Supervision, Conceptualization, Methodology, Validation. Jeroen de Bresser: Investigation, Validation. Laurens Beljaards: Data curation, Software. Jakob Meineke: Resources, Software. Elwin de Weerdt: Conceptualization, Resources. Huangling Lu: Investigation. Mariya Doneva: Supervision, Conceptualization, Methodology, Validation, Writing – review & editing. Marius Staring: Funding acquisition, Project administration, Supervision, Conceptualization, Methodology, Validation, Writing – review & editing.
@@
The authors declare the following financial interests or personal relationships which may be considered as potential competing interests: Chinmay Rao and Laurens Beljaards declare the grant and research support they receive received from Philips. Nicola Pezzotti, Jakob Meineke, Elwin de Weerdt, and Mariya Doneva declare their employment at Philips. Given his role as Associate Editor, Marius Staring had no involvement in the peer review of this article and had no access to information regarding its peer review. Full responsibility for the editorial process for this article was delegated to another journal editor. If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v4

@@ 7 Conclusion
In this work, we introduced PnP-CoSMo, a modular approach to multi-contrast reconstruction which combines content/style modeling with iterative reconstruction, offering the reconstruction quality of state-of-the-art end-to-end methods with improved cross-contrast generalizability despite significantly lenient training data requirements. At its core is the content consistency operation, which provides regularization at the semantic level of the image content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-CoSMo provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In our preliminary radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-CoSMo produced images of at least fairly diagnostic overall quality at =5, enabling 60% more acceleration over clinical reconstructions. To progress from our current proof-of-concept towards a practical implementation, future work will focus mainly on improving and extending the content/style model, reducing reconstruction latency, tackling the problem of model hallucinations, incorporating online registration, and conducting a large-scale radiological study.
@@
Chinmay Rao: Conceptualization, Methodology, Data curation, Software, Investigation, Validation, Formal analysis, Visualization, Writing – original draft. Matthias van Osch: Funding acquisition, Supervision, Conceptualization, Validation, Writing – review & editing. Nicola Pezzotti: Funding acquisition, Supervision, Conceptualization, Methodology, Validation. Jeroen de Bresser: Investigation, Validation. Mark van Buchem: Investigation, Validation. Laurens Beljaards: Data curation, Software. Jakob Meineke: Resources, Software. Elwin de Weerdt: Conceptualization, Resources. Huangling Lu: Investigation. Mariya Doneva: Supervision, Conceptualization, Methodology, Validation, Writing – review & editing. Marius Staring: Funding acquisition, Project administration, Supervision, Conceptualization, Methodology, Validation, Writing – review & editing.
@@
The authors declare the following financial interests or personal relationships which may be considered as potential competing interests: Chinmay Rao and Laurens Beljaards declare the grant and research support they receive from Philips. Nicola Pezzotti, Jakob Meineke, Elwin de Weerdt, and Mariya Doneva declare their employment at Philips. If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.